OpenStudy (anonymous):

Determine which ordered pair in the form (x, y) is a solution of the given equation. (x + 3)y = 14 (11, 1) (3, 2) (5, 2) (7, 2)

jimthompson5910 (jim_thompson5910):

Let's pick on an random answer choice. Say choice D (7,2) means x = 7 and y = 2 ------------------------------------------------------- (x + 3)y = 14 (7 + 3)*2 = 14 ... plug in x = 7 and y = 2 (10)*2 = 14 20 = 14 ... this is false since the last equation is false, this means (7,2) is NOT a solution to (x + 3)y = 14 So we can eliminate choice D